OpenStudy (anonymous):

well i got multiplying decimals down what about dividing decimals like : 0.10 ÷ 2.5 = ... i got 25

Parth (parthkohli):

There are a lot of methods, y'know. Do you know how to divide fractions?

OpenStudy (anonymous):

no i had used a calulator :[ sowwy i didnt learn it

Parth (parthkohli):

Hmm... it's really easy. Do you know how to multiply fractions at least?

OpenStudy (anonymous):

yes i just learned

Parth (parthkohli):

Cool... can you convert 0.10 and 2.5 to fractions?

OpenStudy (anonymous):

can someone explan a easy way to do this

OpenStudy (anonymous):

Method two: Multiply both numbers by 100 (by moving the decimal point of each two places to the right. 0.10 ÷ 2.50 -> 10 ÷ 250 Set up as fraction \[\large \frac{10}{250}\] and reduce, or use long division: 

OpenStudy (anonymous):

ohh okay but how would you divid this sowwy i stupid today

OpenStudy (anonymous):

Parth's method would look like this: \[\large 0.10 ÷ 2.5 = \frac{1}{10} ÷ 2\frac{1}{2}\] \[\large \rightarrow \frac{1}{10} ÷ \frac{5}{2} = \frac{1}{10} × \frac{2}{5}\] \[\large \rightarrow \frac{1}{\cancel{10}5} × \frac{\cancel{2}1}{5} = \frac{1}{25}\] Though, since the question was asked using decimals, you should answer using decimals.

OpenStudy (anonymous):

so then i would covert that 1/25 to a decimal?

OpenStudy (anonymous):

Yep, and you do that the same as I showed above. 

OpenStudy (anonymous):

ohh cool so i used a culculater how would i divid a nmber by a decimal

OpenStudy (anonymous):

If you have a calculator, just enter it in how you see it.

OpenStudy (anonymous):

oh okay i will try

OpenStudy (anonymous):

is it 0.04

OpenStudy (anonymous):

Yep. 1/25 = 0.04